OpenStudy (anonymous):

∫ tan^(-1) (πx)-tan^(-1) (x)dx ,defined from 0 to 2 suggestion: write the integrand as an integral

OpenStudy (anonymous):

\[\int arctan(x)=xarctan(x)-\frac{1}{2}ln(1+x^2)\]

OpenStudy (anonymous):

I do not understand when I say: suggestion: write the integrand as an integral explain me

OpenStudy (anonymous):

i am not sure i do either. arctan is the antiderivative of \[\frac{1}{1+t^2}\]